# Asymmetrical thermonuclear supernovae triggered by the tidal disruption of white dwarfs

### v1

@@
In a dense star cluster core, a tidal disruption event (TDE) of a white dwarf (WD) can occur if the WD passes within the tidal radius of an intermediate-mass black hole (IMBH). Very close encounters cause extreme tidal compression in the WD, raising temperatures enough to induce runaway fusion and produce a thermonuclear supernova (SN). Using the hydrodynamics code AREPO augmented with a 55-isotope nuclear reaction network, we performed high-resolution simulations of the TDE of a $0.6\,\mathrm{M}_{\odot}$ C/O WD by a $500\,\mathrm{M}_{\odot}$ IMBH for different values of the scaled impact parameter $b$ (i.e., the ratio of periapsis distance to tidal radius). Closer encounters produce combined TDE+SN events, with a partial burning of $^{12}\rm{C}$ and $^{16}\rm{O}$ into heavier isotopes – the $^{56}\rm{N}$i fractions of the disrupted WD material vary from 1% at $b=0.19$ to 82% at $b=0.10$, while wider ones ($b\gtrsim 0.20$) lead to standard TDEs. In all cases, the material away from the denser regions remains unburnt, spanning a wide range of radial velocities. Such WD TDEs also exhibit a central cavity, wherein little material is found below a radial velocity of several $1000\,\mathrm{km}\,\mathrm{s}^{-1}$. We also performed 1D and 2D radiative-transfer calculations for these WD-TDEs using the codes CMFGEN and LONG_POL, respectively, covering epochs from a few days to one hundred days. We recover the typical rise times and peak luminosities of SNe Ia, but with an extremely strong viewing-angle dependence of both light curves and spectra. At nebular times, isolated strong emission lines like [Ca ii] $\lambda\lambda$7291, 73237291,\,7323 may appear both displaced and skewed by many $1000\,\mathrm{km}\,\mathrm{s}^{-1}$ – such extreme offsets are harder to identify at earlier times due to optical depth effects and line overlap. WD TDEs may produce a diverse set of transients with extreme asymmetry and peculiar composition.
@@
Stars: white dwarfs – Hydrodynamics – Stars: supernovae: general – Nuclear reactions, nucleosynthesis, abundances – Black hole physics
@@ 2 Hydrodynamics and stellar physics
We employed the 3D hydrodynamics code AREPO (Springel, 2010; Pakmor et al., 2016; Weinberger et al., 2020) to simulate the tidal disruption of a WD by an IMBH. In contrast to the two popular schemes – the Eulerian finite-volume static grid and the Lagrangian smoothed particle hydrodynamics (SPH) – AREPO is an arbitrary Lagrangian-Eulerian moving-mesh code that solves the fluid equations on an unstructured Voronoi grid moving (approximately) with the fluid bulk velocity. It has the advantages of both methods, e.g., the lack of need for artificial viscosity and superior shock treatment of static grid codes, and the inherent adaptive resolution and Lagrangian invariance of SPH codes.
@@ 2.1 White dwarf
Similarly to Pakmor et al. (2012), we constructed the density and pressure profiles of a $0.6\,\mathrm{M}_{\odot}$, nonrotating C/O WD (50% $^{12}\rm{C}$ and 50% $^{16}\rm{O}$) [^1] C/O ratios depend on masses and metallicities of initial stars, and impact future nuclear burning., with central density $\rho_{\mathrm{c,WD}}=3.7\times 10^{6}\,\mathrm{g}\,\mathrm{cm}^{-3}$ and radius $r_{\mathrm{WD}}=8400\,\mathrm{km}$. This was achieved by assuming an initial temperature of $T=5\times 10^{7}\,\mathrm{K}$ and solving for hydrostatic equilibrium. We then mapped these 1D profiles to a 3D AREPO mesh comprised of almost cubical cells of nearly equal masses arranged in HEALPix shells (Górski et al., 2005; Ohlmann et al., 2017). The choice of WD mass was motivated by the fact that the WD mass distribution peaks at $\sim 0.6\,\mathrm{M}_{\odot}$ (e.g., Kepler et al., 2007; Tremblay et al., 2011; Kleinman et al., 2013).
Introducing artificial damping to remove spurious velocities, we then relaxed the WD for ten sound-crossing timescales $t_{\mathrm{cross,WD}}=2\sum_{i=1}^{n}c_{\mathrm{s}}(r_{i})\ (r_{i}-r_{i-1})=12.7\,\mathrm{s}$, where $r_{0}=0$, $r_{n}=r_{\mathrm{WD}}$, and $c_{\mathrm{s}}(r_{i})$ is the speed of sound at radial distance $r_{i}$. We chose a WD cell resolution of $10^{-7}\penalty 10000\ \,\mathrm{M}_{\odot}/\mathrm{cell}$ by performing TDE test runs with different resolution values, namely $10^{-6}\penalty 10000\ \,\mathrm{M}_{\odot}/\mathrm{cell}$, $5\times 10^{-6}\penalty 10000\ \,\mathrm{M}_{\odot}/\mathrm{cell}$ and $10^{-7}\penalty 10000\ \,\mathrm{M}_{\odot}/\mathrm{cell}$. We found that the saturation levels of massive isotope production after periapsis passage, among other quantities, converge for cell resolutions of $5\times 10^{-6}\penalty 10000\ \,\mathrm{M}_{\odot}/\mathrm{cell}$ and $10^{-7}\penalty 10000\ \,\mathrm{M}_{\odot}/\mathrm{cell}$.
@@ 2.2 Black hole
Figure 1: Snapshots of TDEs of $0.6\,\mathrm{M}_{\odot}$ WDs, due to $500\,\mathrm{M}_{\odot}$ IMBHs (white points) for eight different values of impact parameter $b$ at $\sim 500\,\mathrm{s}$. The top two (bottom two) panels show slices of densities and temperatures for $b=0.20,0.19,0.18,0.17$ ($b=0.16,0.15,0.12,0.10$), respectively. As $b$ decreases, the ejecta are more spread out and increasingly unlike a standard TDE, indicating runaway nuclear burning, e.g., the case of $b=0.10$. The masses of the unbound ejecta are shown in Table 1.
@@ 3.2 Nuclear abundances
As expected, deeper encounters lead to further nuclear burning due to increasingly favorable conditions. Consequently, with decreasing b, $^{12}\rm{C}$ and $^{16}\rm{O}$ fractions decrease, intermediate isotope (28Si, 32S, 36Ar, 40Ca, etc.) fractions first increase and then decrease, and $^{56}\rm{N}$i (and other massive isotopes) fraction increase. Such properties are standard of thermonuclear fusion and lead to ejecta with relative abundances between iron-group elements (IGEs), intermediate-mass elements (IMEs), and unburnt C/O that are roughly in line with SN Ia models, whether one considers Chandrasekhar mass or sub-Chandrasekhar mass WDs (see, for example, Seitenzahl et al. 2013). The main differences are the relatively lower abundances of IMEs (either material remains unburnt or tends to burn all the way to IGEs), the wider range of values for the $^{56}\rm{N}$i fraction (from as low as zero at $b=0.20$ up to $\sim 82\%$ at $b=0.10$, which extends beyond the $^{56}\rm{N}$i fraction of 61% in Chandrasekhar mass model DDC0 or 73% of sub-Chandrasekhar mass model SCH7p0 of Blondin et al. 2017). This extreme range thus covers from explosions more typical of stripped-envelope (or perhaps ultra-stripped) SNe of Type Ic (C/O-rich debris for $b=0.19$), to faint SNe Ia with a low $^{56}\rm{N}$i ejected mass (e.g., 91bg-like events; Filippenko et al. 1992a) up to the most IGE-dominated SNe Ia as inferred for 91T-like events (see, e.g., Filippenko et al. 1992b; Schmidt et al. 1994; Phillips et al. 2024).
Figure 3 is a collage of in-plane slices of isotope densities of $^{16}\rm{O}$, 28Si, and $^{56}\rm{N}$i at $\sim 500\,\mathrm{s}$. When $b=0.20$, most of the ejecta are composed of $^{12}\rm{C}$ (not shown) and $^{16}\rm{O}$, and the fractions of $^{56}\rm{N}$i increase with decreasing $b$. The onion-like structure of the ejecta is quite evident, with $^{56}\rm{N}$i, 28Si, and $^{16}\rm{O}$ dominating the inner, intermediate, and outer regions relative to the centers of the debris plumes. This onion-like structure also holds for other isotopes, with more massive isotopes residing in inner shells. For example, $^{12}\rm{C}$, like $^{16}\rm{O}$ is found in the outer regions; 32S, 36Ar, 40Ca, etc., as 28Si, are present in the intermediate regions; and so on. It is important to note that, since the centers of the debris plumes are not at rest but move away from the black hole (unlike standard SNe Ia that are point explosions), this stratification is not present in the radial velocity space; i.e., ejecta at all velocities are composed of material of different compositions, contrary to standard SNe Ia.
@@ 3.3 Ejecta geometry and dynamics
In a standard TDE where the initial orbit of the star is assumed to be parabolic, $\sim 50\%$ of the initial mass is bound to the BH, and $\sim 50\%$ becomes unbound (e.g., Hills, 1988; Rees, 1988). However, extra energy from nuclear fusion results in larger fractions of mass becoming unbound and escaping with large velocities. Table 1 quantifies the dynamics of the unbound ejecta: the masses $m_{\mathrm{unbound}}$, mass fractions $f_{\mathrm{unbound}}$, kinetic energies $E_{\mathrm{K,unbound}}$, and nickel masses $m$ ($^{56}\rm{N}$i). When $b=0.20$, $\sim 56\%$ of the initial WD becomes unbound from the IMBH, similar to a standard TDE. As $b$ decreases, the unbound mass increases, with the $b=0.10$ encounter resulting in $\sim 85\%$ of the initial WD mass becoming unbound. Therefore, the kinetic energies of the ejecta also increase with decreasing $b$, from $\sim 10^{51}\,\mathrm{erg}$ when $b=0.20$ to $\sim 1.6\times 10^{51}\,\mathrm{erg}$ when $b=0.10$. The evolution of the ejecta at later times using 1D radiation hydrodynamics (see Section 4) shows fallback of some of the unbound material to the IMBH. Finally, the nickel masses increase with decreasing $b$, with almost $0.5\,\mathrm{M}_{\odot}$ of $^{56}\rm{N}$i being ejected when $b=0.10$.
Table 1: Masses $m_{\mathrm{unbound}}$, mass fractions $f_{\mathrm{unbound}}$, kinetic energies $E_{\mathrm{K,unbound}}$, and nickel masses $m$ ($^{56}\rm{N}$i) of the unbound ejecta of WD TDEs for different $b$ values ($m_{\mathrm{WD}}=0.6\,\mathrm{M}_{\odot}$). The presented values are averages of the last $20\penalty 10000\ t_{\mathrm{dyn}}$ ($\sim 34\,\mathrm{s}$) of each simulation.
@@
$m_{\mathrm{unbound}}$ $\mathrm{M}_{\odot}$] $0.34$ $0.36$ $0.41$ $0.46$ $0.47$ $0.47$ $0.50$ $0.51$
$f_{\mathrm{unbound}}$ $0.57$ $0.60$ $0.68$ $0.77$ $0.78$ $0.78$ $0.83$ $0.85$
$E_{\mathrm{K,unbound}}$ $10^{51}$ erg] $0.98$ $0.95$ $0.76$ $1.10$ $1.10$ $1.08$ $1.32$ $1.60$
@@ 4 Radiative transfer calculations for SNe Ia from TDEs
Figure 7: Velocity versus ‘spherical’ mass (i.e., $\int 4\pi r^{2}\rho dr$) obtained with the radiation-hydrodynamics code V1D for the $b=0.15$ case at multiple times up to one day (top) and for our full model set at one day (bottom).
@@ 4.1 Preparation of inputs at $1\,\mathrm{d}$
The angular average of the distribution of material along the x-x direction at $\sim 500\,\mathrm{s}$ was remapped into the radiation-hydrodynamics code V1D (Livne, 1993; Dessart et al., 2010). In the process, we averaged the composition within each radial slice, hence losing the heterogeneous composition present in the 3D debris. As shown in Figure 3, unburnt material (i.e., C and O) is located in a ‘shell’ outside the metal-rich, internal regions of the asymmetric ejecta, and thus stands apart from explosive ashes traced by, for example, $^{56}\rm{N}$i. However, because of the peculiar geometry, this unburnt material is present at all ejecta velocities, from the slower-moving regions close to the BH and around the plume all the way out to the largest velocities away from the BH. In contrast, unburnt material in standard, quasispherical (i.e., relative to SNe Ia TDEs) Type Ia SN ejecta survives only in the outermost ejecta layers, and thus only at the largest velocities. This different chemical stratification is a fundamental difference between the ejecta of standard SNe Ia and those from WD-TDEs.
Figure 8: Composition profiles for models $b=0.10$ (top left), $b=0.15$ (top right), and $b=0.19$ (bottom left) as well as the density profiles for the whole model set (bottom right), all at about $1\,\mathrm{d}$ and used as initial conditions for the 1D radiative transfer calculations. We show a few of the dominant species in the ejecta. The Ni profile accounts for radioactive decay (essentially all Ni is due to $^{56}\rm{N}$i).
@@
In contrast to Type Ia SN ejecta, which reach homologous expansion after a few tens of seconds, a SN Ia TDE is influenced by the gravitational pull of the central (here of $500\,\mathrm{M}_{\odot}$) BH, and the ejecta are still far from reaching homologous expansion at the end of our simulations. Consequently, even though the input from AREPO was truncated at small radii to include only unbound material at $\sim 500\,\mathrm{s}$, we found that some material falls back into the BH during the subsequent evolution. As illustrated for the $b=0.15$ case in the top panel of Figure 7, there are two visible effects. First, about a tenth of a solar mass falls back within a few hours after the start of the simulation, which creates a central ‘hole’ in the debris since no material is ejected with velocities below about $3000\,\mathrm{km}\,\mathrm{s}^{-1}$. Second, the velocities of the debris layers above drop by up to a factor of two during this expansion out of the gravitational potential well, reaching homologous expansion ($V\propto r$) only after several hours. Further out in the debris, this pull is weaker, and the material follows a ballistic trajectory earlier. The inner velocity hole varies in size within our model set (bottom panel of Figure 7). It is the largest in the model with the lowest $^{56}\rm{N}$i mass, which also has the lowest kinetic energy and the highest density in the inner debris at $500\,\mathrm{s}$. In other models, the jump varies in size and extent, with no clear monotonic relation to $^{56}\rm{N}$i mass.

### v2

@@
In a dense star cluster core, a tidal disruption event (TDE) of a white dwarf (WD) can occur if the WD passes within the tidal radius of an intermediate-mass black hole (IMBH). Very close encounters cause extreme tidal compression in the WD, raising temperatures enough to induce runaway fusion and produce a thermonuclear supernova (SN). Using the hydrodynamics code AREPO augmented with a 55-isotope nuclear reaction network, we performed high-resolution simulations of the TDE of a $0.6\,\mathrm{M}_{\odot}$ C/O WD by a $500\,\mathrm{M}_{\odot}$ IMBH for different values of the scaled impact parameter $b$ (i.e., the ratio of periapsis distance to tidal radius). Closer encounters produce combined TDE+SN events, with a partial burning of $^{12}\rm{C}$ and $^{16}\rm{O}$ into heavier isotopes – the $^{56}\rm{N}$i fractions of the disrupted WD material vary from 1% at $b=0.19$ to 82% at $b=0.10$, while wider ones ($b\gtrsim 0.20$) lead to standard TDEs. In all cases, the material away from the denser regions remains unburnt, spanning a wide range of radial velocities. Such WD TDEs also exhibit a central cavity, wherein little material is found below a radial velocity of several $1000\,\mathrm{km}\,\mathrm{s}^{-1}$. We also performed 1D and 2D radiative-transfer calculations for these WD-TDEs using the codes CMFGEN and LONG_POL, respectively, covering epochs from a few days to one hundred days. We recover the typical rise times and peak luminosities of SNe Ia, but with an extremely strong viewing-angle dependence of both light curves and spectra. At nebular times, isolated strong emission lines such as [Ca ii] $\lambda\lambda$7291, 73237291,\,7323 may appear both displaced and skewed by many $1000\,\mathrm{km}\,\mathrm{s}^{-1}$ – such extreme offsets are harder to identify at earlier times due to optical depth effects and line overlap. WD TDEs may produce a diverse set of transients with extreme asymmetry and peculiar composition.
@@
Stars: white dwarfs – Hydrodyanamics – Stars: supernovae: general – Nuclear reactions, nucleosynthesis, abundances – Black hole physics
@@ 2 Hydrodynamics and stellar physics
We employed the 3D hydrodynamics code AREPO (Springel, 2010; Pakmor et al., 2016; Weinberger et al., 2020) to simulate the tidal disruption of a WD by an IMBH. In contrast to the two popular schemes – the Eulerian finite-volume static grid and the Lagrangian smoothed particle hydrodynamics (SPH) – AREPO is an arbitrary Lagrangian-Eulerian moving-mesh code that solves the fluid equations on an unstructured Voronoi grid moving (approximately) with the fluid bulk velocity. It has the advantages of both methods; for example, the lack of need for artificial viscosity and superior shock treatment of static grid codes, and the inherent adaptive resolution and Lagrangian invariance of SPH codes.
@@ 2.1 White dwarf
Similarly to Pakmor et al. (2012), we constructed the density and pressure profiles of a $0.6\,\mathrm{M}_{\odot}$, nonrotating C/O WD (50% $^{12}\rm{C}$ and 50% $^{16}\rm{O}$) [^1] C/O ratios depend on masses and metallicities of initial stars, and impact future nuclear burning., with central density $\rho_{\mathrm{c,WD}}=3.7\times 10^{6}\,\mathrm{g}\,\mathrm{cm}^{-3}$ and radius $r_{\mathrm{WD}}=8400\,\mathrm{km}$. This was achieved by assuming an initial temperature of $T=5\times 10^{5}\,\mathrm{K}$ and solving for hydrostatic equilibrium. We then mapped these 1D profiles to a 3D AREPO mesh comprised of almost cubical cells of nearly equal masses arranged in HEALPix shells (Górski et al., 2005; Ohlmann et al., 2017). The choice of WD mass was motivated by the fact that the WD mass distribution peaks at $\sim 0.6\,\mathrm{M}_{\odot}$ (e.g., Kepler et al., 2007; Tremblay et al., 2011; Kleinman et al., 2013).
We set an ambient density of $10^{-10}\,\mathrm{g}\,\mathrm{cm}^{-3}$ for the “vacuum” cells surrounding the WD. These cells are much larger than the WD cells and span a box ten times larger than the farthest distance between the WD and the IMBH. The volume refinement criterion affects the vacuum cells in the immediate vicinity of the WD resulting in a smoothing of its surface, but this effect is relatively negligible. The mass refinement criterion does not apply to these cells.
Introducing artificial damping to remove spurious velocities, we then relaxed the WD for ten sound-crossing timescales $t_{\mathrm{cross,WD}}=2\sum_{i=1}^{n}c_{\mathrm{s}}(r_{i})\ (r_{i}-r_{i-1})=12.7\,\mathrm{s}$, where $r_{0}=0$, $r_{n}=r_{\mathrm{WD}}$, and $c_{\mathrm{s}}(r_{i})$ is the speed of sound at radial distance $r_{i}$. Figure 16 shows the density profiles of the initial and relaxed WDs. Even after $10\,t_{\mathrm{cross,WD}}$, the density profiles remain roughly the same. We chose a WD cell resolution of $10^{-7}\penalty 10000\ \,\mathrm{M}_{\odot}/\mathrm{cell}$ by performing TDE test runs with different resolution values, namely $10^{-6}\penalty 10000\ \,\mathrm{M}_{\odot}/\mathrm{cell}$, $5\times 10^{-6}\penalty 10000\ \,\mathrm{M}_{\odot}/\mathrm{cell}$ and $10^{-7}\penalty 10000\ \,\mathrm{M}_{\odot}/\mathrm{cell}$. We found that the saturation levels of massive isotope production after periapsis passage, among other quantities, converge for cell resolutions of $5\times 10^{-6}\penalty 10000\ \,\mathrm{M}_{\odot}/\mathrm{cell}$ and $10^{-7}\penalty 10000\ \,\mathrm{M}_{\odot}/\mathrm{cell}$.
@@ 3 Results from hydrodynamics
Figure 1: Snapshots of TDEs of $0.6\,\mathrm{M}_{\odot}$ WDs, due to $500\,\mathrm{M}_{\odot}$ IMBHs (white points) for eight different values of impact parameter $b$ at $\sim 500\,\mathrm{s}$. The top two (bottom two) panels show slices of densities and temperatures for $b=0.20,0.19,0.18,0.17$ ($b=0.16,0.15,0.12,0.10$), respectively. As $b$ decreases, the ejecta are more spread out and increasingly unlike a standard TDE, indicating runaway nuclear burning (e.g., the case of $b=0.10$). The masses of the unbound ejecta are shown in Table 1.
@@ 3.2 Nuclear abundances
As expected, deeper encounters lead to further nuclear burning due to increasingly favorable conditions. Consequently, with decreasing b, $^{12}\rm{C}$ and $^{16}\rm{O}$ fractions decrease, intermediate isotope (28Si, 32S, 36Ar, 40Ca, etc.) fractions first increase and then decrease, and $^{56}\rm{N}$i (and other massive isotopes) fraction increase. Such properties are standard of thermonuclear fusion and lead to ejecta with relative abundances between iron-group elements (IGEs), intermediate-mass elements (IMEs), and unburnt C/O that are roughly in line with SN Ia models, whether one considers Chandrasekhar mass or sub-Chandrasekhar mass WDs (see, for example, Seitenzahl et al. 2013). The main differences are the relatively lower abundances of IMEs (either material remains unburnt or tends to burn all the way to IGEs), the wider range of values for the $^{56}\rm{N}$i fraction (from as low as zero at $b=0.20$ up to $\sim 82\%$ at $b=0.10$, which extends beyond the $^{56}\rm{N}$i fraction of 61% in Chandrasekhar mass model DDC0 or 73% of sub-Chandrasekhar mass model SCH7p0 of Blondin et al. 2017). This extreme range thus covers from explosions more typical of stripped-envelope (or perhaps ultra-stripped) SNe of Type Ic (C/O-rich debris for $b=0.19$), to faint SNe Ia with a low $^{56}\rm{N}$i ejected mass (e.g., 91bg-like events; Filippenko et al. 1992a) up to the most IGE-dominated SNe Ia as inferred for 91T-like events (e.g., Filippenko et al. 1992b; Schmidt et al. 1994; Phillips et al. 2024).
Figure 3 is a collage of in-plane slices of isotope densities of $^{16}\rm{O}$, 28Si, and $^{56}\rm{N}$i at $\sim 500\,\mathrm{s}$. When $b=0.20$, most of the ejecta are composed of $^{12}\rm{C}$ (not shown) and $^{16}\rm{O}$, and the fractions of $^{56}\rm{N}$i increase with decreasing $b$. The onion-like structure of the ejecta is quite evident, with $^{56}\rm{N}$i, 28Si, and $^{16}\rm{O}$ dominating the inner, intermediate, and outer regions relative to the centers of the debris plumes. This onion-like structure also holds for other isotopes, with more massive isotopes residing in inner shells. For example, $^{12}\rm{C}$, similar to $^{16}\rm{O}$, is found in the outer regions; 32S, 36Ar, 40Ca, etc., as 28Si, are present in the intermediate regions; and so on. It is important to note that, since the centers of the debris plumes are not at rest but move away from the black hole (unlike standard SNe Ia that are point explosions), this stratification is not present in the radial velocity space; i.e., ejecta at all velocities are composed of material of different compositions, contrary to standard SNe Ia.
@@ 3.3 Ejecta geometry and dynamics
[…]
Table 1: Masses $m_{\mathrm{unbound}}$, mass fractions $f_{\mathrm{unbound}}$, kinetic energies $E_{\mathrm{K,unbound}}$, and nickel masses $m$ ($^{56}\rm{N}$i) of the unbound ejecta of WD TDEs for different $b$ values ($m_{\mathrm{WD}}=0.6\,\mathrm{M}_{\odot}$). [^4] The presented values are averages of $\sim 50$$100\,\mathrm{s}$ after periapsis passage. $E_{\mathrm{K,unbound}}$ is not constant but gradually increases during the course of the simulation.
@@
$m_{\mathrm{unbound}}$ $\mathrm{M}_{\odot}$] $0.28$ $0.31$ $0.38$ $0.43$ $0.44$ $0.45$ $0.48$ $0.47$
$f_{\mathrm{unbound}}$ $0.47$ $0.52$ $0.63$ $0.72$ $0.73$ $0.75$ $0.80$ $0.78$
$E_{\mathrm{K,unbound}}$ $10^{51}$ erg] $0.36$ $0.43$ $0.63$ $0.88$ $1.01$ $1.08$ $1.20$ $1.37$
@@ 4 Radiative transfer calculations for SNe Ia from TDEs
Figure 7: Velocity versus “spherical” mass (i.e., $\int 4\pi r^{2}\rho dr$) obtained with the radiation-hydrodynamics code V1Dfor the $b=0.15$ case at multiple times up to $1\,\mathrm{d}$ (top) and the full model set at $1\,\mathrm{d}$ (bottom). A velocity hole is observed in the unbound debris below $\sim 3000\,\mathrm{km}\,\mathrm{s}^{-1}$ due to fallback of material.
@@ 4.1 Preparation of inputs at $1\,\mathrm{d}$
The angular average of the distribution of material along the x-x direction at $\sim 500\,\mathrm{s}$ was remapped into the radiation-hydrodynamics code V1D (Livne, 1993; Dessart et al., 2010). In the process, we averaged the composition within each radial slice, hence losing the heterogeneous composition present in the 3D debris. As shown in Figure 3, unburnt material (i.e., C and O) is located in a shell outside the metal-rich, internal regions of the asymmetric ejecta, and thus stands apart from explosive ashes traced by, for example, $^{56}\rm{N}$i. However, because of the peculiar geometry, this unburnt material is present at all ejecta velocities, from the slower-moving regions close to the BH and around the plume all the way out to the largest velocities away from the BH. In contrast, unburnt material in standard, quasispherical (i.e., relative to SNe Ia TDEs) Type Ia SN ejecta survives only in the outermost ejecta layers, and thus only at the largest velocities. This different chemical stratification is a fundamental difference between the ejecta of standard SNe Ia and those from WD-TDEs.
Figure 8: Composition profiles for models $b=0.10$ (top left), $b=0.15$ (top right), and $b=0.19$ (bottom left), as well as density profiles for the whole model set (bottom right), all at $\sim 1\,\mathrm{d}$ and used as initial conditions for the 1D radiative transfer calculations. We show a few of the dominant species in the ejecta. The Ni profile accounts for radioactive decay (essentially all Ni is due to $^{56}\rm{N}$i). All elements are present at different radial velocities of the ejecta.
@@
In contrast to Type Ia SN ejecta, which reach homologous expansion after a few tens of seconds, a SN Ia TDE is influenced by the gravitational pull of the central (here of $500\,\mathrm{M}_{\odot}$) BH, and the ejecta are still far from reaching homologous expansion at the end of our simulations. Consequently, even though the input from AREPO was truncated at small radii to include only unbound material at $\sim 500\,\mathrm{s}$, we found that some material falls back into the BH during the subsequent evolution. As illustrated for the $b=0.15$ case in the top panel of Figure 7, there are two visible effects. First, about a tenth of a solar mass falls back within a few hours after the start of the simulation, which creates a central “hole” in the debris since no material is ejected with velocities below about $3000\,\mathrm{km}\,\mathrm{s}^{-1}$. Second, the velocities of the debris layers above drop by up to a factor of two during this expansion out of the gravitational potential well, reaching homologous expansion ($V\propto r$) only after several hours. Further out in the debris, this pull is weaker, and the material follows a ballistic trajectory earlier. The inner velocity hole varies in size within our model set (bottom panel of Figure 7). It is the largest in the model with the lowest $^{56}\rm{N}$i mass, which also has the lowest kinetic energy and the highest density in the inner debris at $500\,\mathrm{s}$. In other models, the jump varies in size and extent, with no clear monotonic relation to $^{56}\rm{N}$i mass.